OpenStudy (anonymous):

Solve. Let logbA=2.627 and logbB=0.348. Find logb AB

OpenStudy (anonymous):

This is one of your log rules.

OpenStudy (anonymous):

\[\log_{b} (x) + \log_{b} (y) = \log_{b} (xy) \]

OpenStudy (anonymous):

So it's 2.975

OpenStudy (anonymous):

Hard to say without knowing b.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I'm not saying you should give up on it. Just asking if you solved for b.

OpenStudy (anonymous):

You know: \(\log_{b} (A) + \log_{b} (B) = \log_{b} (AB)\) \(\log_{b} (A) = 2.627\) \(\log_{b} (B) = 0.348\) So: \(2.627 + 0.348 = \log_{b} (2.627 \times 0.348)\)

OpenStudy (anonymous):

\(2.975 = \log_{b}(.9142)\) \(2.975 = \frac{\log_{10} (.9142)}{\log_{10} (b)}\) \(\log b = \frac{-0.089706}{2.975} = -0.03015\) \(b = 10^{-0.03015} = 0.932925\) Knowing that, you can solve \(\log_{b} (AB)\)